# Learn To be Efficient: Build Structured Sparsity in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success with their billion-level parameters, yet they incur high inference overheads. The emergence of activation sparsity in LLMs provides a natural approach to reduce this cost by involving only parts of the parameters for inference. Existing methods only focus on utilizing this naturally formed activation sparsity, overlooking the potential for further amplifying this inherent sparsity. In this paper, we hypothesize that LLMs can learn to be efficient by achieving more structured activation sparsity. To achieve this, we introduce a novel algorithm, Learn-To-be-Efficient (LTE), designed to train efficiency-aware LLMs to learn to activate fewer neurons and achieve a better trade-off between sparsity and performance. Furthermore, unlike SOTA MoEfication methods, which mainly focus on ReLU-based models, LTE can also be applied to LLMs like GPT and LLaMA with soft activation functions. We evaluate LTE on four models and eleven datasets. The experiments show that LTE achieves a better trade-off between sparsity and task performance. For instance, LTE with LLaMA provides a 1.83x–2.59x FLOPs speed-up on language generation tasks, outperforming the state-of-the-art methods.
@@ 1 Introduction
Large language models (LLMs), like GPT-3 (Brown et al., 2020), OPT (Zhang et al., 2022a), and LLaMA (Touvron et al., 2023a, b), demonstrated emergent natural language ability as model parameters scale to the billion level. Those works have pushed state-of-the-art performance in a wide range of natural language processing applications. However, as the number of model parameters increases, LLMs also become quite expensive for deployment, challenged by the significant computational and memory costs during inference (Liu et al., 2023). Such expensive costs not only significantly inflate the deployment cost but also affect the user experience in many latency-sensitive applications, such as chatbots (Achiam et al., 2023) and autonomous driving (Li et al., 2023a). Various strategies have been introduced to improve LLM inference efficiency, like quantization (Frantar et al., 2022), model pruning (Ma et al., 2023), token sparsity (Zhang et al., 2023; Xiao et al., 2023), and activation sparsity (Liu et al., 2023; Zhang et al., 2022b).
During LLM inference, Feed Forward Network (FFN) layers have been one primary bottleneck, taking over the FLOPs and IOs (Liu et al., 2023). Recent works (Li et al., 2023b; Liu et al., 2023) report that ReLU-based LLMs have extremely high activation sparsity in those FFN layers. For example, in a 175B OPT model, more than 95% activations in FFN layers are zeros (Liu et al., 2023). In light of this favorable balance between accuracy and speedup, existing works explore grouping intermediate neurons in FFN layers into structured chunks (Mirzadeh et al., 2023; Liu et al., 2023), thus achieving hardware-friendly acceleration on modern hardware like GPU (Liu et al., 2023).
Previous efforts focus on utilizing this activation sparsity of ReLU-based LLMs by MoEfication (Zhang et al., 2022b). Given an activation-sparse model, state-of-the-art (SOTA) methods (Zhang et al., 2022b; Liu et al., 2023) propose to convert FFN layers to MoE layers by grouping neurons in the FFN intermediate layer into experts and then applying a router to select most important experts for each input token. Unlike training the MoE model from scratch, which requires a huge amount of resources for pretraining, MoEfication converts FFN layers in pretrained LLMs into MoE layers with relatively small training resources. However, SOTA methods only focus on utilizing this naturally formed activation sparsity of LLMs, overlooking the potential for further amplifying/increasing this inherent sparsity. Moreover, for LLMs with soft activation functions, such as SwiGLU (Shazeer, 2020) in LLaMA, existing works have to replace activation with ReLU to enable sparsity. Unfortunately, our evaluations across many datasets show that this replacement can hurt model performance. Consequently, we need a novel method to improve LLM sparsity under soft activation functions.
In this paper, we hypothesize that LLMs can learn to be efficient and achieve more structured activation sparsity. As shown in Figure 1, our key insight is that, besides optimizing models for task performance alone, we can also train LLMs to be efficiency-aware by developing more structured sparsity, which can be utilized more effectively. However, creating structured sparsity in pretrained LLMs is a non-trivial task due to three practical challenges:
@@
How to group neurons into experts? Since neurons in the same expert are always activated together, a smart grouping strategy that groups functionality-similar neurons is crucial to preserving model performance.
@@
How to train routers more stably? The existing widely used Top-k Softmax routing can lead to a severe accuracy drop (Section3.2). How to jointly train routers and model in a MoEfication setting is still under-explored.
3.
@@
To address these challenges, we introduce Learn-To-be-Efficient (LTE), a novel training algorithm to train efficiency-aware models. LTE integrates an efficiency loss penalty, encouraging models to activate fewer neurons in their FFN layers. Additionally, LTE adopts a threshold-based Sigmoid routing strategy to select experts and employs a two-stage training mechanism to improve the training stability, achieving a more flexible selection of experts instead of selecting a fixed number of experts for all layers and inputs.
To evaluate the effectiveness of our method, we implement LTE on both encoder-based models (RoBERTa, RoBERTa (Liu et al., 2019)) and decoder-based models (GPT2-Medium (Radford et al., 2019), LLaMA-7B (Touvron et al., 2023a)) with the HuggingFace’s transformers, and evaluate our method on eleven datasets from both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks. For NLU tasks, our experiments show that LTE achieves 80% - 95% sparsity without performance drops. Similarly, LTE also achieves a better trade-off between sparsity and task performance than other baselines. For instance, LLaMA-7B with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks, outperforming all other baseline methods. Our evaluation results demonstrate that LTE effectively builds more structured sparsity even on LLMs with soft activation functions.
@@ 2 Related Work
Mixture of Experts (MoE). MoE was proposed by Jacobs et al. a few decades ago to build an integral system with subset networks. Recently, MoE layers have been used in the Transformer architecture as a substitute for the MLP block (Shazeer et al., 2018; Fedus et al., 2022), with the recent Mistral-7B model (Jiang et al., 2023) as a prominent example. In MLP blocks of MoE transformers, instead of using an FFN layer for calculating output, MoE layers construct multiple FFN layers and employ a router to choose a subset of experts to do conditional computation. Even though transformers can have billions of parameters, only a subset of experts are activated for each token, thus effectively increasing model activation sparsity (i.e., the number of skipped neurons in execution).
LLM Sparsity. Recent studies (Li et al., 2023b; Liu et al., 2023) show that trained ReLU-based transformers naturally show a great sparsity in the activation of FFN layers. For example, in a 175B OPT model, more than 95% activations in FFN layers are zeros (Liu et al., 2023). While SOTA models are increasingly using diverse soft activations, like GeLU, SiLU, and SwiGLU, existing work show that replacing those soft activations with ReLU and fine-tuning the model can significantly increase the activation sparsity at the cost of a slight performance drop (Zhang et al., 2022b; Mirzadeh et al., 2023; Li et al., 2023b). However, our evaluation shows that this benefit does not always hold for all datasets (Section 5.2).
Speedup from Sparsity. The emerging sparsity indicates that a large portion of neurons are unnecessary in LLMs inference, which can be taken advantage to improve LLM inference efficiency. One way to utilize sparsity is MoEfication (Zhang et al., 2022b). Recent papers (Liu et al., 2023; Zhang et al., 2022b) show that a single FFN layer in pretrained LLMs can be converted to an MoE layer by splitting the matrices of FFN layers into exclusive sub-matrices and employs a router to activate only neurons in a subset of experts. Similar to MoE models, MoEfication can effectively accelerate LLM inference by only using part of the parameters (Zhang et al., 2022b; Liu et al., 2023).
@@ 3 Background and Motivation
In this section, we first briefly introduce the MoEfication of FFN layers in transformers. Then, we present a study on the limitation of applying noisy top-K Softmax routing to train transformers for better sparsity.
@@ 3.1 Background: MoEfication
FFN layers occupy 60% of the total model computation (Liu et al., 2023) but exhibit high activation sparsity. In contrast to unstructured sparsity methods like weight pruning, activation sparsity in ReLU-based FFN layers is more hardware-friendly, since intermediate neurons in FFN layers can be grouped as structured chunks (Mirzadeh et al., 2023; Liu et al., 2023). MoEfication (Zhang et al., 2022b) is a way to group functional partitions in FFN layers, thereby converting FFN layers to MoE layers for better utilizing the activation sparsity. As known, if a transformer’s hidden states and FFN intermediate dimension are and , respectively, the FF layers process the input as following:
where , are weight metrics, are biased term and is an activation function111SwiGLU in LLaMA needs two matrices to calculate , and the neurons in FFNs can still be used to form experts.. MoEfication aims to group neurons in the FFN intermediate layer into experts. Then, a router is trained for each FFN layer to select most important experts to improve inference sparsity. A recent work, Deja Vu (Liu et al., 2023), also uses a similar setting but treats each neuron as an expert.
However, SOTA MoEfication methods overlook the potential for further optimizing the activation sparsity of LLMs. In this paper, besides optimizing LLMs for task performance, we propose to optimize models for fast inference as well (Figure 1). By designing efficiency-related loss, we find that our novel training method drives LLMs to achieve better sparsity with marginal performance drops. Besides, for LLMs employing soft activations such as GeLU (Hendrycks & Gimpel, 2016) and SwiGLU (Shazeer, 2020), MoEfication (Zhang et al., 2022b) proposes to replace soft activations with ReLU and fine-tune the model to improve activation sparsity. Although recent works (Zhang et al., 2022b; Mirzadeh et al., 2023) show that replacing soft activation functions with ReLU only has a marginal impact on model performance, our evaluations indicate that this conclusion is not universally applicable across all datasets. It also is crucial to develop methods for effectively utilizing sparsity in LLMs with soft activation functions.
@@ 3.2 Limitations of Noisy Top-K Softmax Routing
A potential choice to train the model to adapt to MoE layers is noisy top-K Softmax routing (Fedus et al., 2022; Shazeer et al., 2016), which is a commonly adopted approach to train MoE models from scratch, like Switch Transformers (Fedus et al., 2022). This method entails selecting experts based on the highest router outputs and then calculating Softmax values for these selected outputs. The output of the MoE layer is determined by summing the products of these softmax values with their corresponding experts’ outputs, which is critical for differentiating router parameters. Additionally, a small Gaussian noise is added to the router’s outputs during training to encourage exploration across different experts, which addresses the issue of unselected experts being non-differentiable. However, our empirical findings surprisingly indicate that noisy top-K does not converge well and encounters a biased expert score issue in MoEfication tasks.
To study the effectiveness of noisy top-K Softmax routing, we follow the standard MoEfication setting (Zhang et al., 2022b) to form experts in FFN layers: we fine-tune a model on a downstream dataset and group neurons into experts via parameter clustering. Subsequently, we integrate Softmax routers into models and jointly train both the model and routers with noisy top-K routing. The evaluation results on two datasets are shown in Figure 2. Surprisingly, we observe similar accuracy drops across all sparsity levels, including at very low levels of sparsity. To study the reason behind this, we analyze the expert scores produced by routers and uncover a biased expert issue in MoE layers: one expert in each MoE layer receives an expert score close to 1, while the expert scores for all other routers are nearly 0. This means that only one expert contributes to the inference in each MoE layer, which causes the performance drops.
Before MoEfication, expert outputs in FFN layers can be considered to be multiplied by an expert score equal to 1. The sum of expert scores in Softmax routing is constrained to 1, which results in many experts’ outputs being multiplied by expert scores significantly smaller than 1. Meanwhile, fine-tuning LLMs involves a relatively low learning rate and fewer training iterations. This makes it difficult for neurons in FFN layers to adapt to such largely reduced expert scores. Consequently, the model tends to rapidly converge towards allocating the entire Softmax budget to the most important expert, assigning other experts with expert scores close to 0.
Another limitation of the noisy top-K routing approach is its requirement for a fixed, pre-defined number of selected experts for all inputs and layers. However, the sparsity level can vary depending on the specific inputs and layers. A more adaptive expert selection strategy helps achieve a better trade-off between model efficiency and quality.
@@ 4 Methodology: Learn To be Efficient
In this section, we present, LTE (Learn To be Efficient), to learn more structured MoE in FFN layers for fast inference.
Challenges. Converting a pre-trained LLM’s MLP block into a structured Mixture of Experts (MoE) leads to three practical challenges:
1.
How to group neurons into experts? Since neurons in the same expert are always activated together, A smart grouping strategy that groups functionality-similar neurons is crucial to preserving model performance.
2.
How to train routers more stably? The existing widely used Top-k Softmax routing can lead to a severe accuracy drop (Section3.2). How to jointly train routers and model in a MoEfication setting is still under-explored.
3.
How to select the right experts in serving? The number of experts needed can depend on specific inputs and layers. A more adaptive selection strategy leads to a better trade-off between model efficiency and quality.
@@ 4.1 Experts Grouping
Expert grouping aims to group neurons that are often activated together, thereby improving efficiency by reducing the total number of activated neurons. Here, constructing too many experts (e.g., each neuron as an expert) can significantly increase the cost, while too few can lead to poor model quality. Moreover, we need to group neurons without a significant performance drop on the pre-trained model.
Following the previous work (Zhang et al., 2022b), we use parameter clustering to split MLP neurons into different experts. Fundamentally, in FFN layers, each neuron is associated with a column of (which is a vector). Parameter clustering split first treats the corresponding vector in as the feature for a neuron, then applies the balanced K-Means algorithm (Malinen & Fränti, 2014) to cluster neurons into clusters. Each cluster has neurons by default and will be treated as an individual MoE expert.222 For FFN layers using SwiGLU (e.g., LLaMA), there are two linear networks to calculate neuron values: We use the weight of the gate network, , to cluster neurons.
@@ 4.2 Adaptive Expert Routing
After allocating neurons to different experts, we need to train and orchestrate routers to select experts for input tokens.
4.2.1 Router structure
Similar to existing MoE designs, we use a single fully-connected layer as the router network. The input to the router is the output of the preceding self-attention block, and the router output is the expert score for each expert.
Routing function. To address the biased expert score issue (Section 3.2), instead of using conventional Softmax routing, we employ the Sigmoid function on router outputs to get expert scores:
where is the input to the corresponding FFN layer, is the index for the expert, and is the router network weight for the -th expert. Unlike the Softmax function, the Sigmoid function computes each expert score independently, which circumvents the biased expert score due to the constraint on the sum of outputs in the Softmax function.
Threshold-based experts selection. SOTA methods (Zhang et al., 2022b; Fedus et al., 2022) select a fixed number of experts for all layers and inputs. However, the necessary number of experts may differ depending on inputs and layers. Here, we propose to select experts more adaptively for different inputs and layers for a better trade-off between model sparsity and quality. Specifically, we leverage a threshold-based method to enable adaptive selection:
where is the output for the -th expert, is the indicator function, and is a predefined threshold. Only the experts whose score is larger than will be activated. Threshold-based routing does not have to select a fixed number of experts in each MoE layer. Instead, it selects experts with sufficiently large expert scores. Consequently, within the same MoE layer, as the router generates larger expert scores, more experts are selected.

### v3

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success with their billion-level parameters, yet they incur high inference overheads. The emergence of activation sparsity in LLMs provides a natural approach to reduce this cost by involving only parts of the parameters for inference. However, existing methods only focus on utilizing this naturally formed activation sparsity in a post-training setting, overlooking the potential for further amplifying this inherent sparsity. In this paper, we hypothesize that LLMs can learn to be efficient by achieving more structured activation sparsity. To achieve this, we introduce a novel training algorithm, Learn-To-be-Efficient (LTE), designed to train efficiency-aware LLMs to learn to activate fewer neurons and achieve a better trade-off between sparsity and performance. Furthermore, unlike SOTA MoEfication methods, which mainly focus on ReLU-based models, LTE can also be applied to LLMs like LLaMA using non-ReLU activations. Extensive evaluation on language understanding, language generation, and instruction tuning tasks show that LTE consistently outperforms SOTA baselines. Along with our hardware-aware custom kernel implementation, LTE reduces LLaMA2-7B inference latency by 25% at 50% sparsity.
@@ 1 Introduction
Large language models (LLMs), like GPT-3 [2], OPT [42], and LLaMA [35, 36], have demonstrated impressive natural language ability. However, the skyrocketing number of model parameters has not only inflated the deployment costs of LLMs, due to their significant computational and memory demands during inference [20], but also affected the user experience in many latency-sensitive applications, such as chatbots [1] and autonomous driving [16]. Recent advances have been leveraging sparsity to improve LLM inference efficiency, including model weight quantization [8] and pruning [21], token sparsity [44, 41], and activation sparsity [20, 43].
During LLM inference, Feed Forward Network (FFN) layers are the primary efficiency bottleneck, accounting for over 60% of the FLOPs and I/O operations [20], but they exhibit high activation sparsity, especially in ReLU-based LLMs [17]. For example, in a 175B OPT model, more than 95% of activations in FFN layers are zeros [20]. Recent advances leverage this activation sparsity via MoEfication [43, 24]: they convert FFN layers to MoE layers by grouping neurons in the FFN intermediate layer into experts and then applying a router to select most important experts for each input token. Unlike training the MoE model from scratch, MoEfication converts FFN layers in pretrained LLMs into MoE layers with relatively small training resources.
However, existing advances only focus on manipulating the pre-trained activation sparsity of neurons, overlooking the potential for further increasing this inherent sparsity. Moreover, they are limited to the historically ReLU-based LLMs, whereas emerging advanced models use soft activation functions for better model quality (e.g., SwiGLU in LLaMA [29] and GeGLU in Gemma [34]), exhibiting much lower natural activation sparsity. While existing works suggest replacing the model activation with ReLU to enable sparsity [43, 24], we show that this can hurt model performance (Section 5.2).
In this paper, we hypothesize that LLMs can learn to be efficient and achieve more structured activation sparsity. As shown in Figure 1, our key insight is that, without compromising model quality, we can also train LLMs to be efficiency-aware by developing more structured sparsity, which is more friendly for achieving hardware speedup. However, creating structured sparsity in pretrained LLMs is non-trivial due to two practical challenges:
@@
How to train routers more stably? The widely-used Top-k Softmax routing can lead to a severe accuracy drop (Section3.2). How to jointly train the model and routers in a MoEfication setting is still open-ended.
@@
To address these challenges, we introduce Learn-To-be-Efficient (LTE), a novel training algorithm to train efficiency-aware models. LTE integrates an efficiency loss penalty, encouraging models to activate fewer neurons in their FFN layers while keeping good task performance. Additionally, LTE adopts a threshold-based Sigmoid routing strategy to select experts and employs a two-stage training mechanism to improve training stability. LTE achieves a more flexible selection of experts instead of selecting a fixed number of experts for all layers and inputs. Same as Deja Vu [20] and moefication [43], LTE provides very structured sparsity. We further develop a custom CUDA kernel with Triton, a Python-like CUDA programming language, to enable wall-clock time speedup from such structured sparsity (Section 4.3).
We evaluate LTE on both encoder-based models (RoBERTa, RoBERTa [19]) and decoder-based models (GPT2-Medium [27], LLaMA2-7B [35]) with the HuggingFace’s transformers. Our extensive experiments on Natural Language Understanding (NLU) tasks, downstream Natural Language Generation (NLG) tasks, and instruction tuning tasks, show that LTE consistently outperforms state-of-the-art designs. For instance, LLaMA with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks without compromising model quality. After integrating our hardware-efficient implementation of sparse matrix multiplication kernel, LTE reduces LLaMA2-7B 25% wall-clock time latency at around 50% sparsity. Our evaluation results demonstrate that LTE effectively builds more structured sparsity even on LLMs with soft activation functions.
@@ 2 Related Work
Mixture of Experts (MoE). MoE was proposed by [13] a few decades ago to build an integral system with subset networks. Recently, MoE layers have been used in the Transformer architecture as a substitute for the MLP block [30, 6], with the recent Mistral-7B model [14] as a prominent example. In MLP blocks of MoE transformers, instead of using an FFN layer for calculating output, MoE layers construct multiple smaller FFN layers and employ a router to choose a subset of experts to do conditional computation. Even though transformers can have billions of parameters, only a subset of experts are activated for each token, thus effectively increasing model activation sparsity (i.e., the number of skipped neurons in execution).
LLM Contextual Sparsity. Recent studies [17, 20] show that trained ReLU-based transformers naturally show a great sparsity in the activation of FFN layers. For example, in a 175B OPT model, more than 95% activations in FFN layers are zeros [20]. While SOTA models are increasingly using diverse soft activations, like GeLU and SwiGLU, existing work shows that replacing soft activations with ReLU and fine-tuning the model can increase the activation sparsity without greatly hurting model quality [43, 24, 17]. Yet, we find that this benefit is not consistent for all datasets (Section 5.2). This emerging sparsity indicates that a large portion of neurons are unnecessary in LLM inference (i.e., sparsity), which we can leverage to improve LLM inference efficiency like using MoEfication [43]. Recent papers [20, 43] show that a single FFN layer in pretrained LLMs can be converted to an MoE layer, by splitting the matrices of FFN layers into exclusive sub-matrices and employing a router to activate only neurons in a subset of experts. Similar to MoE models, MoEfication can effectively accelerate LLM inference by only using part of the parameters [43, 20].
Model Weight Sparsity. Another orthogonal direction regarding model sparsity is static model weight sparsity [7, 33], which often prunes the model weights and keeps them static during inference. For instance, Wanda [33] estimates model weight importance and prunes unimportant weights, so all inputs will use the same subset of weights. In contrast, contextual sparse models select a different subset of weights for different inputs. Nonetheless, model pruning can be applied to contextual sparse models too, meaning a complementary optimization to contextual sparsity.
@@ 3 Background and Motivation
In this section, we first briefly introduce the MoEfication of FFN layers in transformers. Then, we present a study on the limitation of applying noisy top-K Softmax routing in the Moefication setting.
@@ 3.1 Background: MoEfication
MoEfication [43] is a way to group neurons in FFN layers, thereby converting FFN layers to MoE layers for better execution speedup. For a transformer whose hidden states and FFN intermediate dimension are and , respectively, the FFN layers process the input as follows:
where , are weight metrics, are biased term and is an activation function. MoEfication aims to group neurons in the FFN intermediate layer into experts. Then, a router is trained for each FFN layer to select experts (k < n), thus reducing activation load, to speed up inference. A recent work, Deja Vu [20], uses a similar setting but treats each neuron as an expert.
However, SOTA MoEfication methods overlook the potential for further optimizing the activation sparsity of LLMs. Besides, for LLMs employing soft activations such as GeLU [11] and SwiGLU [29], MoEfication [43] proposes to replace soft activations with ReLU and fine-tune the model to improve activation sparsity. Although recent works [43, 24] show that this replacement has a marginal impact on model performance, we notice that this hypothesis is not widely applicable (Section 5.2).
@@ 3.2 Limitations of Noisy Top-K Softmax Routing
One potential solution to train and convert the model into MoE layers is noisy top-K Softmax routing [6, 31]. It selects experts based on the highest router outputs and then calculates Softmax values for these selected outputs. The output of the MoE layer is determined by summing the products of these softmax values with their corresponding experts’ outputs. Additionally, a small Gaussian noise is added to the router’s outputs during training to encourage exploration across different experts, which addresses the issue of unselected experts being non-differentiable.
To study the effectiveness of noisy top-K Softmax routing, we follow the standard MoEfication setting [43] to form experts in FFN layers: we fine-tune a model on a downstream dataset and group neurons into experts via parameter clustering. Subsequently, we integrate Softmax routers into models and jointly train both the model and routers. As shown in Figure 2, we observe noticeable accuracy drops even at very low sparsity levels. When diving into the expert scores produced by routers, we observe a very biased expert score allocation: only one expert in an MoE layer has an expert score close to 1, while other experts receive nearly 0 scores. This results in only one expert contributing to the inference in each MoE layer, leading to performance drops.
A potential reason behind this biased allocation is that the sum of expert scores in Softmax routing is constrained to 1. Unlike traditional MoE models, which typically select 1-2 experts [30, 6], MoEfied models need to select a much larger number of experts [20, 43], but the sum of experts score is still 1; this can complicate the allocation of the Softmax budget and cause a biased allocation.
@@ 4 Methodology: Learn To be Efficient
In this section, we present LTE (Learn To be Efficient) to learn more structured MoE in FFN layers for fast inference. Specifically, we tackle two key challenges toward practical MoEfication for diverse LLMs: (1) How to train routers more stably? (Section 4.2); and (2) How to select the right experts in serving? (Section 4.2.2).
@@ 4.1 Experts Grouping
The first step of LTE is to group neurons in FFN layers to form experts. Expert grouping aims to group neurons that are often activated together, thereby improving efficiency by reducing the total number of activated neurons. Here, constructing too many experts (e.g., each neuron as an expert) can significantly increase the cost, while too few can lead to poor model quality. Moreover, we need to group neurons without a significant performance drop on the pre-trained model.
Following the previous work [43], we group 32 neurons as an expert in FFN layers, and use parameter clustering to group MLP neurons into different experts. Fundamentally, in FFN layers, each neuron is associated with a column of (which is a vector). Parameter clustering first treats the corresponding vector in as the feature for a neuron, then applies the balanced K-Means algorithm [22] to cluster neurons into clusters. Each cluster has neurons by default and will be treated as an individual MoE expert.111 For FFN layers using SwiGLU (e.g., LLaMA), there are two linear networks to calculate neuron values: We use the weight of the gate network, , to cluster neurons. We also conduct an ablation study on other alternative grouping strategies in Section 5.4.
@@ 4.2 Adaptive Expert Routing
4.2.1 Router Design
After constructing the experts, we need to decide on the right routing strategy. Similar to existing MoE designs, we use a fully-connected layer as the router network. The input to the router is the output of the preceding self-attention block, and the router output is the expert score for each expert.
Routing function. To address the biased expert score issue (Section 3.2), we employ the Sigmoid function on router outputs to get expert scores. Unlike the Softmax function, the Sigmoid function computes each expert score independently, which circumvents the biased expert score due to the constraint on the sum of outputs in the Softmax function:
where is the input to the corresponding FFN layer, is the index for the expert, and is the router network weight for the -th expert.
Threshold-based experts selection. SOTA methods [43, 6] select a fixed number of experts for all layers and inputs. However, the necessary number of experts may differ depending on inputs and layers. Here, we propose to select experts more adaptively for different inputs and layers for a better trade-off between model sparsity and quality, by leveraging a threshold-based method to enable adaptive selection. The MoEfied FFN layers become:
where is the output for the -th expert, is the indicator function, and is a predefined threshold. Only experts with score larger than will be activated. Consequently, within the same MoE layer, as the router generates larger expert scores, more experts are selected.